OpenStudy (anonymous):

4x-9y=6.5 7y-2x=-4.5

OpenStudy (angela210793):

@Jerzey u haven't understood how this kind of system is solved yet? (ask wht u don't understand)

jhonyy9 (jhonyy9):

multiplied with every two equation you get 8x-18y=13 8x-18y=13 - assumed -4x+14y=-9 /*2 --- -8x+28y=-18 -------------- 0 10y=-5 --- y=-5/10 = -1/2 - so than y=-1/2 --- 8x=13+18(-1/2) 8x=13-18/2 8x=13-9 8x=4 x=4/8 x=1/2 x=1/2 y=-1/2